140 mg of KN3 is dissolved into 3.5l of water what is the ph of the solution?

I got 2.76. I think i may need to subtract the value from 14 to get the true pH or is 2.76 the true pH?

Any details on the matter are greatly appreciated.

This is all the information given in the question.

I did concentration = moles/vol

then used the concentration to find the pH with pH=-log[concentration]. Is that Wrong?
Answer #1

we have

140 mg of KN3 is dissolved into 3.5l of water

mole of KN3 = mass / molar mass = 0.140 g / 81.1184 g/mole = 1.726 * 10^-3 mole.

and

concentration of KN3 =  1.726 * 10^-3 mole / 3.5 L= 4.931 * 10^-4 M

pKa of HN3 is 4.63.

Here

salt hydrolysis occurs.

It is the salt of strong base and weak acid.

Using

pH = 1/2 * [PKw + pKa + log C]

or

pH = 1/2 * [14 + 4.63 + log (4.931 * 10^-4 )]

or

pH = 7.66